Probability and Statistics I GSIA, Carnegie Mellon University

45-733, Spring 2002

Homework #3, due Tuesday, February 6, 2002

Please show all relevant work! Please do the following problems from the book:

pp 192-4 3,15
pp 204-7 25,33
pp 214-6 35,43
pp 217-21 47

If you get stuck, it would be wise to try to work even-numbered problems similar to the ones assigned, since many of these have solutions in the back of the book.